### v1

# Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu

Abstract
While large language models (LLMs) like ChatGPT have shown impressive capabilities in Natural Language Processing (NLP) tasks, a systematic investigation of their potential in this field remains largely unexplored. This study aims to address this gap by exploring the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP? To answer these questions, we take the first step to provide a comprehensive overview of LLMs in NLP. Specifically, we first introduce a unified taxonomy including (1) parameter-frozen application and (2) parameter-tuning application to offer a unified perspective for understanding the current progress of LLMs in NLP. Furthermore, we summarize the new frontiers and the associated challenges, aiming to inspire further groundbreaking advancements. We hope this work offers valuable insights into the potential and limitations of LLMs in NLP, while also serving as a practical guide for building effective LLMs in NLP.
Large Language Models Meet NLP: A Survey
Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu Central South University   Harbin Institute of Technology   University of Hong Kong Tsinghua University   University of Illinons at Chicago lbqin@csu.edu.cn, {qgchen,car}@ir.hit.edu.cn
1 Introduction
Recently, large language models (LLMs) represent a significant breakthrough in AI through scaling up language models Zhao et al. (2023a); Kaddour et al. (2023); Yang et al. ; Hadi et al. (2023); Zhuang et al. (2023). Current studies on LLMs, such as GPT-series Brown et al. (2020); Ouyang et al. (2022), PaLM-series Chowdhery et al. (2022), OPT Zhang et al. (2022a), and LLaMA Touvron et al. (2023), have shown impressive zero-shot performance. In addition, LLMs also bring some emergent abilities including instruction following Wei et al. (2022a), chain-of-thought reasoning Wei et al. (2022c) and in-context learning Min et al. (2022), which attract increasing attention Wei et al. (2022b).
With the advancement of large language models, as shown in Figure 1, LLMs allow various natural language processing (NLP) tasks (e.g., zero-shot mathematical reasoning, text summarization, machine translation, information extraction and sentiment analysis) to be achieved through a unified generative paradigm, which has achieved remarkable success Wei et al. (2022c, 2023a); Qin et al. (2023a); Wang et al. (2023a, d, h, j); Wan et al. (2023b); Peng et al. (2023); Huang et al. (2023a). Additionally, some LLMs in NLP work without needing any additional training data and can even surpass traditional models fine-tuned with supervised learning. This advancement significantly contributes to the development of NLP literature. As a result, the community has witnessed an exponential growth of LLMs for NLP studies, which motivates us to investigate the following questions: (1) How are LLMs currently applied to NLP tasks in the literature? (2) Have traditional NLP tasks already been solved with LLMs? (3) What is the future of the LLMs for NLP?
Figure 1: The example of applying LLMs for NLP tasks (e.g., mathematical reasoning, machine translation, information extraction and sentiment analysis).
Figure 2: The taxonomy of LLMs for NLP, including parameter-frozen (a) and parameter-tuning paradigm (b), where blue module with ice denotes that the parameters are kept unchanged, and orange module with fire represents the fine-tuning of full or selected parameters.
To answer the above questions, we make the first attempt to present a comprehensive and detailed analysis on LLMs for NLP. The overarching goal of this work is to explore the current developments in LLMs for NLP. To this end, in this paper, we first introduce the relevant background and preliminary. Furthermore, we introduce a unified paradigm on LLMs for NLP: (1) parameter-frozen application including (i) zero-shot learning and (ii) few-shot learning; (2) parameter-tuning application containing (i) full-parameter tuning and (ii) parameter-efficient tuning, aiming to provide a unified perspective to understand the current progress of LLMs for NLP:
  • Parameter-frozen application directly applies prompting approach on LLM for NLP tasks without the need for parameter tuning. This category includes zero-shot and few-shot learning, depending on whether the few-shot demonstrations is required.
  • Parameter-tuning application refers to the need for tuning parameters of LLMs for NLP tasks. This category includes both full-parameter and parameter-efficient tuning, depending on whether fine-tuning is required for all model parameters.
Finally, we conclude by identifying potential frontier areas for future research, along with the associated challenges to stimulate further exploration.
In summary, this work offers the following contributions:
  • (1)
    First survey: We present the first comprehensive survey of Large Language Models (LLMs) for Natural Language Processing (NLP) tasks.
  • (2)
    New taxonomy: We introduce a new taxonomy including (1) parameter-frozen application and (2) parameter-tuning application, which provides a unified view to understand LLMs for NLP tasks.
  • (3)
    New frontiers: We discuss emerging areas of research in LLMs for NLP and highlight the challenges associated with them, aiming to inspire future breakthroughs.
  • (4)
    Abundant resources: We create the first curated collection of LLM resources for NLP, including open-source implementations, relevant corpora, and a list of research papers. These resources are available at https://github.com/LightChen233/Awesome-LLM-for-NLP.
We expect this work will be a valuable resource for researchers and spur further advancements in the field of LLM-based NLP.
2 Background
[…]
2.1 Parameter-Frozen Paradigm
Parameter-frozen paradigm can directly apply prompting for NLP tasks without any parameter tuning. As shown in Figure 2 (a), this category encompasses zero-shot learning and few-shot learning Brown et al. (2020); Kojima et al. (2022).
Zero-shot Learning
In zero-shot learning, LLMs leverage the instruction following capabilities to solve NLP tasks based on a given instruction prompt, which is defined as:
$\mathcal{P}=\texttt{Prompt}(\mathcal{I}),$ (1)
where $\mathcal{I}$ and $\mathcal{P}$ denote the input and output of prompting, respectively.
Few-shot Learning
Few-shot learning uses in-context learning capabilities to solve the NLP tasks imitating few-shot demonstrations. Formally, given some demonstrations $\mathcal{E}$, the process of few-shot learning is defined as:
$\mathcal{P}=\texttt{Prompt}(\mathcal{E},\mathcal{I}).$ (2)
2.2 Parameter-Tuning Paradigm
As shown in Figure 2 (b), the parameter-tuning paradigm involves adjusting LLM parameters for NLP tasks, covering both full-parameter and parameter-efficient tuning.
Full-parameter Tuning
In the full-parameter tuning approach, all parameters of the model $\mathcal{M}$ are fine-tuned on the training dataset $\mathcal{D}$:
^$\hat{\mathcal{M}}=\texttt{Fine-tune}(\mathcal{M}|\mathcal{D}),$ (3)
where $\hat{\mathcal{M}}$ is the fine-tuned model with the updated parameters.
Parameter-efficient Tuning
Parameter-efficient tuning (PET) involves adjusting a set of existing parameters or incorporating additional tunable parameters (like Bottleneck Adapter Houlsby et al. (2019), Low-Rank Adaptation (LoRA) Hu et al. (2021), Prefix-tuning Li and Liang (2021a), and QLoRA Dettmers et al. (2023)) to efficiently adapt models for specific NLP tasks. Formally, parameter-efficient tuning first tunes a set of parameters $\mathcal{W}$, denoting as:
$\hat{\mathcal{W}}=\texttt{Fine-tune}(\mathcal{W}|\mathcal{D},\mathcal{M}),$ (4)
where $\hat{\mathcal{W}}$ stands for the trained parameters.
3 Natural Language Understanding
As shown in Figure 3, we first describe some typical NLP understanding tasks, which consists of Semantic Analysis (§$\S$3.1), Information Extraction (§$\S$3.2), Dialogue Understanding (§$\S$3.3), and Table Understanding (§$\S$3.4).
3.1 Sentiment Analysis
Sentiment analysis, a key function in natural language processing, identifies the emotional tone of a text, like positive opinions or criticisms Wankhade et al. (2022).
3.1.1 Parameter-Frozen Paradigm
Zero-shot Learning
With the help of instruction tuning, LLMs have been equipped with excellent zero-shot learning ability Belkhir and Sadat (2023). Recent studies Zhang et al. (2023g) find that using simple instructions can elicit ChatGPT’s strong capabilities on a series of sentiment analysis tasks such as sentiment classification and aspect-based sentiment analysis. Moreover, current mainstream LLMs Koto et al. (2024) possess the ability of multilingual understanding to analyze the sentiment conveyed by different languages based on sentiment lexicons Koto et al. (2024).
Few-shot Learning
Few-shot prompting not only elicits in-context learning in LLMs but also elaborates the intent of users more clearly. According to the findings presented by previous studies Zhang et al. (2023g); Zhao et al. (2023b); Xu et al. (2023c), incorporating exemplars to the prompts significantly boosts LLMs’ performance on aspect-based sentiment analysis and emotion recognition tasks. Furthermore, Sun et al. (2023b) introduce few-shot learning on more complex procedures, incorporating multi-LLM negotiation framework for sentiment analysis.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Frozen Paradigm Taxonomy,ver [  Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Zhang et al. (2023g), Koto et al. (2024), Zhao et al. (2023b), Xu et al. (2023c), Sun et al. (2023b) , leaf, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Zhang et al. (2023c),Wei et al. (2023a),Xie et al. (2023), Li and Zhang (2023),Li et al. (2023e), Bi et al. (2023) , leaf, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Pan et al. (2023), He and Garner (2023), Hudeček and Dušek (2023), Heck et al. (2023), Gao et al. (2023a),
Li et al. (2022b), Zhang et al. (2023i, h), Wu et al. (2023c), Das et al. (2023a), Chi et al. (2023), Hu et al. (2022b),
King and Flanigan (2023), Addlesee et al. (2023),Chung et al. (2023),Lee et al. (2023),Lin et al. (2023), Cao (2023) , leaf, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Singha et al. (2023), Patnaik et al. (2024), Ye et al. (2023, 2024), Sui et al. (2023a, b), Cheng et al. (2022),
Zhang et al. (2023f, k, j, 2024), Chen (2023), Luo et al. (2023b), Li et al. (2023b), Jiang et al. (2023) , leaf, text width=48.3em] ] ] [      Generation (§4),ver [Summarization (§4.1) [e.g., Goyal et al. (2022), Ravaut et al. (2023b), Bhaskar et al. (2022), Wang et al. (2023b), Zhang et al. (2023e, b),
Adams et al. (2023), Tang et al. (2023) , leaf, text width=48.3em] ] [Code Generation (§4.2) [e.g., Chen et al. (2021), Nijkamp et al. (2022), Christopoulou et al. (2022), Luo et al. (2023c), Allal et al. (2023),
Li et al. (2023f, g), Guo et al. (2024) Roziere et al. (2023), Zheng et al. (2023), , leaf, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Wei et al. (2023b),Zhu et al. (2023a),Li et al. (2023a, c),Alves et al. (2023),Raunak et al. (2023),Lu et al. (2023b) , leaf, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g., Wei et al. (2022c), Zhang et al. (2022b), Kojima et al. (2022), Wang et al. (2023g), Touvron et al. (2023),
Lu et al. (2023d), Gao et al. (2023b) , leaf, text width=48.3em] ]] ]
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.4em,font=,, where level=2text width=11.3em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [Parameter-Tuning Paradigm Taxonomy,ver [  Understanding (§3),ver [Sentiment Analysis (§3.1) [e.g., Wang et al. (2022), Varia et al. (2022), Yang and Li (2023), Zhao et al. (2016), Qiu et al. (2023) , leaf2, text width=48.3em] ] [Information Extraction
 (§3.2) [e.g., Lu et al. (2023a),Gan et al. (2023),Sainz et al. (2023),Wang et al. (2023f), Das et al. (2023b),Liang et al. (2023) , leaf2, text width=48.3em] ] [Dialogue Understanding
 (§3.3) [e.g., Xie et al. (2022),Zhao et al. (2022a), Gupta et al. (2022),Yu et al. (2022),Feng et al. (2023b), Liu et al. (2023b) , leaf2, text width=48.3em] ] [Table Understanding (§3.4) [e.g., Li et al. (2023d), Xie et al. (2022), Xue et al. (2023), Zhang et al. (2023a), Zhu et al. (2024), Bai et al. (2023),
Zhang et al. (2023d) , leaf2, text width=48.3em] ]] [       Generation (§4),ver [Summarization (§4.1) [e.g., Pagnoni et al. (2022), Zhao et al. (2022b), Yuan et al. (2022), Feng et al. (2023a), Li and Liang (2021b),
Ravaut et al. (2023a) , leaf2, text width=48.3em] ] [Code Generation (§4.2) [e.g., Wang et al. (2021, 2023i),Le et al. (2022),Shojaee et al. (2023),Ayupov and Chirkova (2022),Zhuo et al. (2024),
Weyssow et al. (2023) , leaf2, text width=48.3em] ] [Machine Translation (§4.3) [e.g., Xu et al. (2023b, 2024), Iyer et al. (2023),Moslem et al. (2023),Ustun and Stickland (2022),Alves et al. (2023)
Wu et al. (2023a, 2024), , leaf2, text width=48.3em] ] [Mathematical Reasoning
 (§4.4) [e.g.,  Luo et al. (2023a), Yue et al. (2023), Ho et al. (2023), Schick et al. (2023), Hu et al. (2022a, 2023b) , leaf2, text width=48.3em] ] ] ]
Figure 3: Taxonomy of LLMs for NLP including Parameter-Frozen Paradigm and Parameter-Tuning Paradigm.
3.1.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-parameter instruction tuning has been shown to be an effective approach to bridge the gap between task-agnostic pre-training and task-specific inference. Specifically, Wang et al. (2022) design unified sentiment instruction for various aspect-based sentiment analysis tasks to elicit the LLMs. Varia et al. (2022) utilize task-specific sentiment instructions to fine-tune LLMs for the inter-task dependency. Yang and Li (2023) transform the visual input into plain text during prompt construction for instruction tuning. These works demonstrate the potential of tuning LLMs for advanced sentiment analysis.
Parameter-Efficient Tuning
Sentiment analysis techniques have numerous real-world applications such as opinion mining Zhao et al. (2016). Therefore, efficiency is a vital dimension for evaluating sentiment analysis methods. Qiu et al. (2023) utilize LoRA to tune LLMs on the empathy multi-turn conversation dataset namely SMILECHAT to develop emotional support systems.
3.2 Information Extraction
Information Extraction (IE) tasks aim at extracting structural information from plain text, which typically includes relation extraction (RE), named entity recognition (NER), and event extraction (EE) Xu et al. (2023a).
3.2.1 Parameter-Frozen Paradigm
Zero-shot Learning
Inspired by the impressive capabilities of LLMs on various tasks, recent studies Zhang et al. (2023c); Wei et al. (2023a) begin to explore zero-shot prompting methods to solve IE tasks by leveraging knowledge embedded in LLMs. Wei et al. (2023a) , Xie et al. (2023) and Zhang et al. (2023c) propose a series of methods to decompose question-answering tasks by breaking down NER into smaller, simpler subproblems, which improves the overall process. In addition, Xie et al. (2023) further introduce two methods, syntactic prompting and tool augmentation, to improve LLMs’ performance by incorporating the syntactic information.
Few-shot Learning
Considering the gap between sequence labeling and text generation, providing exemplars could help LLMs better understand the given task and follow the problem-solving steps. To select pertinent demonstrations, Li and Zhang (2023) deploy the retrieval module to retrieve the most suitable examples for the given test sentence. Instead of using natural language for structured output, Li et al. (2023e) and Bi et al. (2023) propose reformulating IE tasks as code with code-related LLMs such as Codex.
3.2.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
A common practice to customize LLMs is fine-tuning LLMs on the collected dataset. There typically are three tuning paradigms adopted to enhance LLMs’ abilities. The first one is tuning LLMs on a single dataset to strengthen a specific ability. The second one is standardizing data formats across all IE subtasks, thus enabling a single model to efficiently handle diverse tasks (Lu et al., 2023a; Gan et al., 2023). The last one is tuning LLMs on a mixed dataset and testing on the unseen tasks Sainz et al. (2023); Wang et al. (2023f), which is always used to improve the generalization ability of LLMs.
Parameter-Efficient Tuning
Tuning huge parameters of LLMs poses a significant challenge to both research and development. To address this challenge, Das et al. (2023b) propose a method for dynamic sparse fine-tuning that focuses on a specific subset of parameters during the IE training process. This approach is particularly useful when dealing with limited data. Meanwhile, Liang et al. (2023) introduce Lottery Prompt Tuning (LPT), a method that efficiently tunes only a portion of the prompt vectors used for lifelong information extraction. This technique optimizes both parameter efficiency and deployment efficiency.
3.3 Dialogue Understanding
Dialogue understanding typically consists of spoken language understanding (SLU) Tur and De Mori (2011); Qin et al. (2019, 2021) and dialogue state tracking (DST) Sarikaya et al. (2016); Jacqmin et al. (2022).
3.3.1 Parameter-Frozen Paradigm
Zero-shot Learning
Recent studies highlight the effectiveness of LLMs in dialogue understanding through zero-shot prompting Pan et al. (2023); He and Garner (2023); Hudeček and Dušek (2023); Heck et al. (2023). Gao et al. (2023a) and Addlesee et al. (2023) introduce zero-shot chain-of-thought prompting strategies in LLMs, enhancing understanding by step-by-step reasoning. Moreover, Zhang et al. (2023i) and Wu et al. (2023c) treat SLU and DST as agent systems and code generation tasks to effectively improve task performance. Further, Chung et al. (2023), Chi et al. (2023) and Zhang et al. (2023h) extend the task to actual scenarios and understand the dialog by zero-shot prompting for efficient interaction and dialog management.
Few-shot Learning
Limited by the instruction following ability of the LLMs, recent studies have focused on improving model performance in dialogue understanding through the relevant few-shot demonstrations Hudeček and Dušek (2023). To address “overfitting” in the given few-shot demonstrations, Hu et al. (2022b), King and Flanigan (2023), Das et al. (2023a), Li et al. (2022b), Lee et al. (2023), King and Flanigan (2023) and Addlesee et al. (2023) further introduce some methods for retrieving diverse few-shot demonstrations to improve understanding performance. Lin et al. (2023) and Cao (2023) integrate DST tasks with an agent through in-context-learning, enhancing dialogue understanding capabilities.
3.3.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-parameter tuning involves not freezing any parameters and using all parameters to train dialogue understanding tasks Yu et al. (2022). Specifically, Xie et al. (2022); Zhao et al. (2022a) unifies structured tasks into a textual format by training full parameters demonstrating significant improvement and generalization. Gupta et al. (2022) utilize input with some demonstrations as a new DST representation format to train LLM with full parameters and achieve great results.
Parameter-Efficient Tuning
Limited by the huge cost of full-parameter fine-tuning, a lot of work begins to focus more on Parameter-Efficient Tuning (PET) for lower-cost dialogue understanding task training. Specifically, Feng et al. (2023b) present LDST, a LLaMA-driven DST framework that leverages LoRA technology for parameter-efficient fine-tuning, achieving performance comparable to ChatGPT. Liu et al. (2023b) provide a key-value pair soft-prompt pool, selecting soft-prompts from the prompting pool based on the conversation history for better PET.
3.4 Table Understanding
Table understanding involves the comprehension and analysis of structured data presented in tables, focusing on interpreting and extracting meaningful information, like Table Question Answering Jin et al. (2022).
3.4.1 Parameter-Frozen Paradigm
Zero-shot Learning
Recently, the advancements for LLMs have paved the way for exploring zero-shot learning capabilities in understanding and interpreting tabular data Singha et al. (2023); Patnaik et al. (2024); Ye et al. (2024). Ye et al. (2023) and Sui et al. (2023a) concentrate on breaking down large tables into smaller segments to reduce irrelevant data interference during table understanding. Further, Patnaik et al. (2024) introduce CABINET, a framework that includes a module for generating parsing statements to emphasize the data related to a given question. Sui et al. (2023b) develop TAP4LLM, enhancing LLMs’ table understanding abilities by incorporating reliable information from external knowledge sources into prompts. Additionally, Ye et al. (2024) propose a DataFrameQA framework to utilize secure Pandas queries to address issues of data leakage in table understanding. These efforts signify a significant stride towards leveraging LLMs for more effective and efficient zero-shot learning in table data comprehension.
Few-shot Learning
Few-shot learning has been an increasingly focal point for researchers to address the limitations of LLMs, particularly in the context of table understanding and instruction following ability Chen (2023); Zhang et al. (2024). Luo et al. (2023b) propose a hybrid prompt strategy coupled with a retrieval-of-thought to further improve the example quality for table understanding tasks. Cheng et al. (2022) introduce Binder to redefine the table understanding task as a coding task, enabling the execution of code to derive answers directly from tables. Furthermore, Li et al. (2023b), Jiang et al. (2023) and Zhang et al. (2023k, f) conceptualize the table understanding as a more complex agent task, which utilizes external tools to augment LLMs in table tasks. Building upon these developments, ReAcTable Zhang et al. (2023j) integrates additional actions into the process, such as generating SQL queries, producing Python code, and directly answering questions, thereby further enriching the few-shot learning landscape for LLMs.
3.4.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Leveraging the existing capabilities of LLMs, Full-Parameter Tuning optimizes these models for specific table understanding tasks. Li et al. (2023d) and Xie et al. (2022) adapt a substantial volume of table-related data for table instruction tuning, which leads to better generalization in table understanding tasks. Additionally, Xue et al. (2023) introduce DB-GPT to enhance LLMs by fine-tuning them and integrating a retrieval-augmented generation component to better support table understanding.
Parameter-Efficient Tuning
Xie et al. (2022) utilize prompt-tuning for efficient fine-tuning within a unified framework of table representation instructions. Moreover, Zhang et al. (2023a), Zhu et al. (2024) and Bai et al. (2023) adapt Low-Rank Adaptation (LoRA) during instruction-tuning for better table understanding and further table cleaning. Furthermore, Zhang et al. (2023d) address challenges related to long table inputs by implementing LongLoRA, demonstrating its efficacy in managing long-context issues in table understanding tasks.
4 Natural Language Generation
This section presents the LLMs for classific NLP generation tasks containing Summarization (§$\S$4.1), Code Generation (§$\S$4.2), Machine Translation (§$\S$4.3), and Mathematical Reasoning (§$\S$4.4), which are illustrated in Figure 3.
4.1 Summarization
Summarization aims to distill the most essential information from a text document, producing a concise and coherent synopsis that retains the original content’s primary themes (Shi et al., 2018).
4.1.1 Parameter-Frozen Paradigm
Zero-shot Learning
In the exploration of zero-shot learning for text summarization, LLMs such as GPT-3 have demonstrated amazing and superior performance in generating concise and factually accurate summaries, challenging the need for traditional fine-tuning approaches (Goyal et al., 2022; Bhaskar et al., 2022; Wang et al., 2023b). Zhang et al. (2023e) highlight instruction tuning as pivotal for LLMs’ summarization success. Ravaut et al. (2023b) scrutinize LLMs’ context utilization, identifying a bias towards initial document segments in summarization tasks. These studies collectively underscore the versatility and challenges of deploying LLMs in zero-shot summarization.
Few-shot Learning
For few-shot learning, LLMs like ChatGPT are scrutinized for their summarization abilities. Zhang et al. (2023b) and Tang et al. (2023) demonstrate that leveraging in-context learning and a dialog-like approach can enhance LLMs’ extractive summarization, particularly in achieving summary faithfulness. Adams et al. (2023) introduce a “Chain of Density” prompting technique, revealing a preference for denser, entity-rich summaries over sparser ones. Together, these studies reveal the evolving strategies to optimize LLMs for summarization tasks.
4.1.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-Parameter Tuning for text summarization leverages the power of LLMs, optimizing them for specific summarization tasks. DIONYSUS (Li et al., 2022a) adapts to new domains through a novel pre-training strategy tailored for dialogue summarization. Socratic Pretraining (Pagnoni et al., 2022) introduces a question-driven approach to improve the summarization process. This allows the model to be easily adapted for different summarization tasks, resulting in more controllable and relevant summaries.
Parameter-Efficient Tuning
PET strategies have revolutionized the adaptability of large pre-trained models for specific summarization tasks, demonstrating the power of fine-tuning with minimal parameter adjustments Feng et al. (2023a). Zhao et al. (2022b) and Yuan et al. (2022) adapt prefix-tuning (Li and Liang, 2021b) for dialogue summarization, enhancing model knowledge and generalization across domains. Ravaut et al. (2023a) develop PromptSum to combine prompt tuning with discrete entity prompts for controllable abstractive summarization. These approaches collectively show the efficacy of PET in enabling robust, domain-adaptive, and controllable summarization with minimal additional computational costs.
4.2 Code Generation
Code generation involves the automatic creation of executable code from natural language specifications, facilitating a more intuitive interface for programming (Chen et al., 2021).
4.2.1 Parameter-Frozen Paradigm
Zero-shot Learning
Recent advancements in code generation have been significantly propelled by the development of LLMs, with studies showcasing their proficiency in generating code in a zero-shot manner. Code LLMs, trained on both code and natural language, have a robust and amazing zero-shot learning capability for programming tasks (Nijkamp et al., 2022; Roziere et al., 2023). Moreover, CodeT5+ enriches the landscape by proposing a flexible encoder-decoder architecture and a suite of pretraining objectives, leading to notable improvements (Wang et al., 2023i). These models collectively push the boundary of what is achievable in code generation, offering promising avenues for zero-shot learning.
Few-shot Learning
Code generation is being revolutionized by few-shot learning. This technique allows models to create precise code snippets by learning from just minimal examples Lu et al. (2021). Chen et al. (2021), Allal et al. (2023), Li et al. (2023f), Luo et al. (2023c) and Christopoulou et al. (2022) illustrate the efficacy of few-shot learning, demonstrating an adeptness at code generation that surpasses its predecessors. The development of smaller, yet powerful models (Li et al., 2023g; Guo et al., 2024), further highlights accessibility of few-shot code generation technologies, making them indispensable tools in the arsenal of modern developers.
4.2.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-parameter tuning represents a pivotal strategy in enhancing code generation models, allowing comprehensive model optimization. Specifically, CodeT series (Wang et al., 2021, 2023i) epitomize this approach by incorporating code-specific pre-training tasks and architecture flexibility, respectively, to excel in both code understanding and generation. CodeRL (Le et al., 2022) and PPOCoder (Shojaee et al., 2023) introduce deep reinforcement learning, leveraging compiler feedback and execution-based strategies for model refinement, whereas StepCoder (Shojaee et al., 2023) advances this further by employing reinforcement learning, curriculum learning and fine-grained optimization techniques. These models collectively demonstrate significant improvements across a spectrum of code-related tasks, embodying the evolution of AI-driven programming aids.
Parameter-Efficient Tuning
PET emerges as a pivotal adaptation in code tasks, striking a balance between performance and computational efficiency (Weyssow et al., 2023). Studies (Ayupov and Chirkova, 2022; Zhuo et al., 2024) exploring adapters and LoRA showcase PET’s viability on code understanding and generation tasks, albeit with limitations in generative performance.
4.3 Machine Translation
Machine translation is a classical task that utilize computers to automatically translate the given information from one language to another, striving for accuracy and preserving the semantic essence of the original material (Bahdanau et al., 2014).
4.3.1 Parameter-Frozen Paradigm
Zero-shot Learning
In the realm of zero-shot learning, Zhu et al. (2023a) and Wei et al. (2023b) enhance LLMs’ multilingual performance through cross-lingual and multilingual instruction-tuning, significantly improving translation tasks. OpenBA contributes to the bilingual model space, demonstrating superior performance in Chinese-oriented tasks with a novel architecture (Li et al., 2023c). These advancements highlight the potential of LLMs in aligning language in zero-shot settings.
Few-shot Learning
In the exploration of few-shot learning for machine translation (MT), recent studies present innovative strategies to enhance the capabilities of LLMs Li et al. (2023a); Huang et al. (2024). Lu et al. (2023b) introduce Chain-of-Dictionary Prompting (CoD) to improve the MT of rare words by in-context-learning in low-resource languages. Raunak et al. (2023) investigate the impact of demonstration attributes on in-context learning, revealing the critical role of output text distribution in translation quality. Together, these works illustrate the significant potential of few-shot learning and in-context strategies in advancing the field of MT with LLMs.
4.3.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-parameter tuning in machine translation with LLMs represents a frontier for enhancing translation accuracy and adaptability Xu et al. (2023b). Iyer et al. (2023) demonstrate the potential of LLMs in disambiguating polysemous words through in-context learning and fine-tuning on ambiguous datasets, achieving superior performance in multiple languages. Moslem et al. (2023) and Wu et al. (2024) focus on exploring fine-tuning methods that enhance real-time and context-aware translation capabilities. Xu et al. (2024) propose Contrastive Preference Optimization (CPO) to refine translation quality further, pushing LLMs towards better performance. These studies reveal the efficacy and necessity of fine-tuning approaches in realizing the full potential of LLMs for complex machine translation tasks.
Parameter-Efficient Tuning
PET is emerging as a transformative approach for integrating LLMs into machine translation (MT), balancing performance and efficiency. Ustun and Stickland (2022) empirically assess PET’s efficacy across different languages and model sizes, highlighting adapters’ effectiveness with adequate parameter budgets. Alves et al. (2023) optimize the finetuning process with adapters, striking a balance between few-shot learning and finetuning efficiency. These studies collectively underline PET’s potential to revolutionize MT by making LLMs more adaptable and resource-efficient.
4.4 Mathematical Reasoning
Mathematical reasoning tasks in NLP involve the use of NLP techniques to understand information from mathematical text, perform logical reasoning, and generate answers Lu et al. (2023e).
4.4.1 Parameter-Frozen Paradigm
Zero-shot Learning
Mathematics serves as a testbed to investigate the reasoning capabilities of LLMs OpenAI (2023); Touvron et al. (2023). The vanilla prompting method asks LLMs to directly arrive at the final answer to a given mathematical problem. It is very challenging and the reasoning process is not transparent to humans. To address it,  Kojima et al. (2022) develop a zero-shot chain-of-thought technique, which utilizes the simple prompt “Let’s think step by step” to elicit mathematical reasoning in LLMs. By doing this, the LLM can break down the problem into smaller, easier-to-solve pieces before arriving at a final answer. Further, Wang et al. (2023g) propose a new decoding strategy, called self-consistency. This approach integrates a series of prompting results to boost the mathematical performance.
Few-shot Learning
Recent studies explore constructing more suitable exemplars for LLMs to improve mathematical reasoning.  Wei et al. (2022c) introduce chain-of-thought prompting, which presents a few chain-of-thought demonstrations to teach LLMs to think step by step. However, manually constructing the demonstrations in few-shot learning is time- and labor-consuming. To solve this problem,  Zhang et al. (2022b) and  Lu et al. (2023d) propose to select in-context examples automatically. Even given detailed examples, it is still hard for LLMs to calculate the numbers precisely. To address this issue, PAL Gao et al. (2023b) directly generates programs as intermediate reasoning steps. These programs are then executed using a runtime environment, like a Python interpreter, to find the better and robust solution.
Figure 4: The future work and new frontier for LLM in NLP tasks.
4.4.2 Parameter-Tuning Paradigm
Full-Parameter Tuning
Full-parameter tuning is a common way to specify LLMs’ behaviors on mathematical reasoning tasks.  Luo et al. (2023a) apply their proposed Reinforcement Learning from Evol-Instruct Feedback (RLEIF) method to the domain of math to improve the mathematical reasoning abilities of LLMs.  Yue et al. (2023) introduce the MathInstruct dataset to enhance the general math problem-solving ability of LLMs through in-domain instruction tuning.  Ho et al. (2023) teach the small language models to perform mathematical reasoning by distilling the generated intermediate rationales by large language models.  Schick et al. (2023) present ToolFormer, which can use the calculator to perform simple numeric calculations when solving math problems.
Parameter-Efficient Tuning
Fine-tuning LLMs with full parameter updates incurs significant memory overhead, limiting accessibility for many users. Parameter-efficient tuning techniques, such as LoRA Hu et al. (2022a), offer a promising alternative. Additionally, Hu et al. (2023b) propose a user-friendly framework for integrating various adapters into LLMs, enabling them to tackle tasks like mathematical reasoning.
Takeaways
(1) LLMs offer a unified generative solution paradigm for various NLP tasks. (2) LLMs in NLP tasks still have a certain gap from smaller supervised learning models. (3) Continuing to fine-tune LLMs on NLP tasks bring substantial improvements.
5 Future Work and New Frontier
In this section, as shown in Figure 4, we highlight some new frontiers, hoping to spur more breakthroughs in the future.
5.1 Multilingual LLMs for NLP
Despite the significant success of LLMs in English NLP tasks, there are over 7,000 languages worldwide. How to extend the success of English-centric LLMs to NLP tasks in other languages is an important research question Qin et al. (2024). Inspired by this, recent research has increasingly focused on using multilingual LLMs to solve NLP tasks in multilingual scenarios Xue et al. (2021); Workshop et al. (2022); Shi et al. (2022); Qin et al. (2023a); Winata et al. (2023).
Two main challenges in this direction are as follows: (1) Enhancing Low-Resource Language Performance: Due to poor performance in low-resource languages, how to build universal multilingual LLMs that achieve promising performance in NLP tasks across languages is a direction worth exploring. (2) Improving Cross-lingual Alignment: The key to multilingual LLMs is improving the alignment between English and other languages. Effectively achieving cross-lingual alignment in cross-lingual NLP tasks is a challenge.
5.2 Multi-modal LLMs for NLP
The current LLMs achieve excellent performance in text modality. However, integrating more modalities is one of the key ways to achieve artificial general intelligence (AGI). Therefore, a lot of work has begun to explore multi-modal LLMs for multi-modal NLP tasks Lu et al. (2022, 2023c); Yang et al. (2023a, b); Zhang et al. (2023l).
The primary challenges in this field are: (1) Complex Multi-modal Reasoning: Currently, most multi-modal LLMs focus on simple multi-modal reasoning, like recognition Wang et al. (2023e); Liu et al. (2023a), while neglecting complex multi-modal reansoning Yang et al. (2023b); Lu et al. (2023c). Therefore, how to effectively explore complex multi-modal reasoning for NLP is a crucial topic. (2) Effective Multi-modal Interaction: Existing methods often simply focus on adding direct multi-modal projection or prompting to LLM for bridge multi-modality gap Wang et al. (2023e); Liu et al. (2023a); Wu et al. (2023b); Mitra et al. (2023). Crafting a more effective multi-modal interaction mechanism in multi-modal LLMs to solve NLP tasks is an essential problem.
5.3 Tool-usage in LLMs for NLP
While LLMs have shown success in NLP tasks, they can still face challenges when applied in real-world scenarios Qin et al. (2023b). Therefore, a lot of work explores utilizing LLMs as central controllers to enable the usage or construction of tools and agents to solve more practical NLP tasks Shinn et al. (2023); Wang et al. (2023c); Zhu et al. (2023b); Hu et al. (2023a).
The primary concerns are: (1) Appropriate Tool Usage: Current works always consider static tool usage, neglecting to choose appropriate tools to use. Identifying the correct tools and using them accurately is a key issue in solving NLP tasks efficiently. (2) Efficient Tool Planning: Current works still focus on the usage of a single tool for NLP tasks. Motivated by this, there is a pressing need for NLP tasks to achieve an efficient tool chain that leverages multiple tools in a coordinated manner. For example, when facing Task-oriented Dialogue tasks, we can use three tools: booking flight tickets, booking train tickets, and booking bus tickets. Then, how to collaborate to make the trip time as short as possible and the cost as low as possible is a typical problem in effective tool planning.
5.4 X-of-thought in LLMs for NLP
When LLMs solve complex NLP problems, they often cannot directly give correct answers and require complex thinking. Therefore, some works adapt X-of-thought (XoT) for advanced logical reasoning. XoT primarily aims to refine logical processing for better NLP task solution Kojima et al. (2022); Zhang et al. (2022b); Qin et al. (2023a); Yao et al. (2023); Chen et al. (2022); Lei et al. (2023).
Key challenges in this direction include: (1) Universal Step Decomposition: How to develop a method for universally applicable step decomposition to generalize LLMs to various NLP tasks is the core challenge of XoT. (2) Prompting Knowledge Integration: Diverse promptings enhance model performance across various scenarios. How to better integrate the knowledge of different XoT to solve NLP problems is an important direction.
5.5 Hallucination in LLMs for NLP
During solving the NLP tasks, LLMs inevitably suffer from the hallucinations where LLMs produce outputs that deviate from world knowledge Muhlgay et al. (2023); Min et al. (2023), user request Adlakha et al. (2023), or self-generated context Liu et al. (2022). This deviation harms the reliability of LLMs in practical scenarios.
The primary challenges in hallucination are: (1) Efficient Hallucination Evaluation: How to find appropriate and unified evaluation benchmarks and metrics for LLMs in various NLP tasks is a key challenge. (2) Leveraging Hallucinations for Creativity: Hallucinations can often stimulate certain creative abilities. How to leverage hallucination to stimulate creativity and generate better innovative knowledge is an interesting topic.
5.6 Safety in LLMs for NLP
Applying large models to downstream NLP tasks also raises inevitable safety concerns, including copyright issues Chang et al. (2023), hate toxicity Hartvigsen et al. (2022), social bias Wan et al. (2023a); Dhamala et al. (2021) and psychological safety Huang et al. (2023b). Inspired by this, a series of works focus on the research on the safety of LLMs for diverse NLP tasks Ganguli et al. (2022); Sun et al. (2023a).
The main challenges to safety in LLMs are: (1) Safety Benchmark Construction: Currently, there are few security-related benchmarks for LLM on various NLP tasks. Establishing effective safety benchmarks is a critical objective in this area. (2) Multilingual Safety Risks: LLM suffers more safety risks across languages and cultures. Identifying and mitigating these risks in a multilingual context is a significant challenge.
6 Conclusion
In this work, we make the first attempt to offer a systemic overview of LLMs in NLP, introducing a unified taxonomy about parameter-frozen applications and parameter-tuning applications. Besides, we highlight new research frontiers and challenges, hoping to facilitate future research. Additionally, we maintain a publicly available resource website to track the latest developments in the literature. We hope this work can provide valuable insights and resources to build effective LLMs in NLP.